OpenStudy (anonymous):

I Will Medal And Fan Best Answer! I need explanation AND answer! Add. 3/4 + 3/10 Write your answer as a fraction in simplest form.

OpenStudy (anonymous):

21

OpenStudy (anonymous):

21?

OpenStudy (anonymous):

just kidding im just messing around

OpenStudy (anonymous):

-_-

OpenStudy (ccswims):

you must find the common denominator between 10 and 4

OpenStudy (anonymous):

i know what to do, just not good at knowing how

OpenStudy (ccswims):

ok, the common denominator is 20, because 10 can go into 20, and 4 can go into 20

OpenStudy (ccswims):

but you cant just leave it as 3/20. for 3/4 you multiplied 5*4 to get 20. so now you must multiply 3*5 also

OpenStudy (ccswims):

what you do to the denominator you must ALWAYS do to the numerator

OpenStudy (ccswims):

so now you have 15/20, correct? because 3/5 = 15 and 4/5 = 20

OpenStudy (anonymous):

i don't get that last part

OpenStudy (ccswims):

now you must do the same to 3/10, except this time you multiplied it by two because 10*2 = 20. so now, instead of multiplying 3*5, you are going to multiply 3*2

OpenStudy (ccswims):

ok, so you need help more on the last part? ok, so when you go to find the common denominator, you have to multiply the same number you multiplied on the bottom, to the top.

OpenStudy (ccswims):

so, for example, if i wanted the common denominator to be 10, and we had the fraction 2/5, you would multiply 5*2 to get 10 right?

OpenStudy (ccswims):

well, you must also multiply 2 to the numerator (2*2) because you did that to the denominator

OpenStudy (ccswims):

get it?

OpenStudy (ccswims):

so basically, whatever you do to the bottom you must also do to the top.

OpenStudy (ccswims):

ok, so now do what we did with 3/4 (and remember we got 15/20) now do it with 3/10

OpenStudy (anonymous):

i don't get most of that........

OpenStudy (anonymous):

Sorry, i'm stupid

OpenStudy (ccswims):

your not stupid, this certain part just may not be your strong suit, for instance my weak part in math is angle proofs and parabolas and postulates. you'll get it tho. tell me where you exactly start to drop. where exactly do you not understand?

OpenStudy (anonymous):

I don't understand MOST math. This is when i lost: so now you have 15/20, correct? because 3/5 = 15 and 4/5 = 20

OpenStudy (ccswims):

you arent dividing. its not 3/5, its 3 * 5. ok, what doesnt make sense to you during that part? why i multiplied the number outcome? Like, do you know why i multiplied 3 * 5 and 4 * 5?

OpenStudy (anonymous):

no

OpenStudy (ccswims):

ok, we need to get the common denominator because you cannot currently add 3/4 + 3/10 because the denominators are different

OpenStudy (ccswims):

20 is the least common denominator we can use.

OpenStudy (ccswims):

do you know why i picked 20?

OpenStudy (anonymous):

no

OpenStudy (ccswims):

ok, i picked 20 because both 4 and 10 can go into 20, right?

OpenStudy (anonymous):

yes

OpenStudy (ccswims):

ok, how do we multiply 4 to 20

OpenStudy (ccswims):

what number multiplies with 4 to get 20?

OpenStudy (anonymous):

5